### v1

# Batch-ICL: Effective, Efficient, and Order-Agnostic In-Context Learning

Abstract
In this paper, by treating in-context learning (ICL) as a meta-optimization process, we explain why LLMs are sensitive to the order of ICL examples. This understanding leads us to the development of Batch-ICL, an effective, efficient, and order-agnostic inference algorithm for ICL. Differing from the standard N-shot learning approach, Batch-ICL employs $N$ separate 1-shot forward computations and aggregates the resulting meta-gradients. These aggregated meta-gradients are then applied to a zero-shot learning to generate the final prediction. This batch processing approach renders the LLM agnostic to the order of ICL examples. Through extensive experiments and analysis, we demonstrate that Batch-ICL consistently outperforms most permutations of example sequences. In some cases, it even exceeds the performance of the optimal order for standard ICL, all while reducing the computational resources required. Furthermore, we develop a novel variant of Batch-ICL featuring multiple “epochs” of meta-optimization. This variant implicitly explores permutations of ICL examples, further enhancing ICL performance.
1 Introduction
[…]
Numerous studies (Olsson et al., 2022; Wang et al., 2023; Dai et al., 2023; Li et al., 2023; Akyürek et al., 2023; Von Oswald et al., 2023; Ren and Liu, 2023; Xie et al., 2022) have contributed to understanding the mechanism of ICL. Specifically, some research (Dai et al., 2023; Ren and Liu, 2023) describes ICL as a meta-optimization where an LLM is utilized as a meta-optimizer. Meta-gradients are produced through forward processing with ICL examples. These meta-gradients are then applied to the language model through the attention mechanism, resulting in an effective ICL model.
We propose that these insights shed light on a commonly recognized issue in LLMs: their performance is highly influenced by the order of ICL examples. As emphasized by (Lu et al., 2022), changing the order of ICL examples can lead to significantly different outcomes. This paper offers a preliminary explanation for this phenomenon. In an $N$-shot ICL process, the meta-gradient is formed based on $N$ sequentially presented examples. Due to the causal attention mechanism in LLMs, the meta-gradient shaped by any given example is indirectly affected by the ones that came before it. Consequently, the order of these examples plays a critical role in shaping the final ICL model. This process is similar to training a neural network with $N$ samples where the batch size is one. In such cases, the gradients generated by each sample are influenced not only by the sample itself but also by the parameter updates from preceding samples, leading to variations in model parameters and overall performance. Given that the gradient from each sample represents a local optimum and is susceptible to causing suboptimal results, we suggest there is much potential for improving standard ICL performance by placing emphasis on the example order.
In this paper, we introduce Batch-ICL, an effective, efficient and order-agnostic inference algorithm for in-context learning. Our method diverges from the standard ICL, which typically employs a single $N$-shot process for an $N$-shot ICL task. Instead, we implement $N$ separate 1-shot forward computations. This is then followed by the aggregation of $N$ corresponding meta-gradients at specific layers. These aggregated meta-gradients are subsequently employed in the forward computation when an LLM handling a zero-shot task, ultimately producing the final predictions.
Batch-ICL addresses concerns related to the order of examples in ICL. Across a variety of tasks, Batch-ICL consistently exhibits improved accuracy compared to the average accuracy achieved through all permutations of ICL examples and occasionally outperforms even the best order permutation. Meanwhile, Batch-ICL reduces the computational resources needed for executing an ICL sequence.
Additionally, we have discovered that although the standard ICL exhibits instability when presented with different ordered sequences, there are advantages to be gained from the interaction among sequential examples. We expand Batch-ICL into a “multi-epoch” variant, which implicitly enumerates the order permutations in a much more efficient manner, leading to further enhancements in performance.
The remainder of this paper is structured as follows: Section 2 offers an introduction to in-context learning viewed from a meta-optimization perspective. Section 3 presents our proposed Batch-ICL, including its algorithm pipeline, a complexity analysis, and multi-epoch extension. We demonstrate the experimental results and analysis in Section 4 and conduct extensive method analysis in Section 5. We discuss related works in Section 6 and conclude our paper in Section 7.
2 Understanding In-Context Learning from a Meta-Optimization Perspective
Many works (Aizerman et al., 1964; Irie et al., 2022; Dai et al., 2023; Ren and Liu, 2023) have demonstrated the similarity between the linear attention and the linear layers optimized by gradient descent. In this section, we briefly review this similarity and then introduce the in-context learning from a meta-optimization perspective.
2.1 Dual Form between Linear Attention and Gradient Descent in Linear Layers
Consider a linear layer defined as:
$f(\boldsymbol{x})=\boldsymbol{W}_{0}\boldsymbol{x},$
where $\boldsymbol{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ represents the initial weight matrix. Given a sequence of input vectors $\boldsymbol{x}_{i}\in\mathbb{R}^{d_{in}},i\in[1,N]$, and their corresponding labels $\boldsymbol{y}_{i}\in\mathbb{R}^{d_{out}}$, the error signal $\boldsymbol{e}_{i}$ is produced by backpropagation, where $\boldsymbol{e}_{i}=-\eta\frac{\partial\mathcal{L}}{\partial y_{i}}$, with $\eta$ as the learning rate and $\mathcal{L}$ as the loss function. The weight matrix updates as follows:
$\boldsymbol{W}^{\prime}=\boldsymbol{W}_{0}+\Delta\boldsymbol{W}=\boldsymbol{W}_{0}+\sum^{N}_{i}\boldsymbol{e}_{i}\otimes\boldsymbol{x}_{i}.$ (1)
Recap that a linear attention is formulated as:
$\texttt{LA}(\boldsymbol{V},\boldsymbol{K},\boldsymbol{q})=\boldsymbol{V}^{\top}\boldsymbol{K}\boldsymbol{q}=\sum_{i}\boldsymbol{v}_{i}(\boldsymbol{k}^{\top}_{i}\boldsymbol{q}).$ (2)
When focusing on the current input $\boldsymbol{x}_{N}$, we can derive the dual form between the linear layer optimized by gradient descent and the linear attention:
$f(\boldsymbol{x}_{N})$ $=(\boldsymbol{W}_{0}+\Delta\boldsymbol{W})\boldsymbol{x}_{N}$ (3)
$=\boldsymbol{W}_{0}\boldsymbol{x}_{N}+\sum^{N-1}_{i}(\boldsymbol{e}_{i}\otimes\boldsymbol{x}_{i})\boldsymbol{x}_{N}$
$=\boldsymbol{W}_{0}\boldsymbol{x}_{N}+\sum^{N-1}_{i}\boldsymbol{e}_{i}(\boldsymbol{x}^{\top}_{i}\boldsymbol{x}_{N})$
$=\boldsymbol{W}_{0}\boldsymbol{x}_{N}+\texttt{LA}(\boldsymbol{E},\boldsymbol{X}_{1:N-1},\boldsymbol{x}_{N}).$
2.2 ICL is an Meta-Optimization
In an $N$-shot ICL setting, a Transformer (Vaswani et al., 2017)-based LLM consists of $K$ layers. The LLM processes an input sequence in the form of $I_{1},l_{1},I_{2},l_{2},\dots,I_{N},l_{N},I_{q}$, where $I_{i}$ and $l_{i}$ together represent a demonstration example, with $I_{i}$ being the input and $l_{i}$ as the corresponding label. Here, $I_{q}$ denotes the genuine query, and the LLM’s task is to predict its label. We represent the embedding of the input sequence at $k$-layer as $\boldsymbol{X}^{k}=[\boldsymbol{x}^{k}_{1},\boldsymbol{x}^{k}_{2},\dots,\boldsymbol{x}^{k}_{q}]$ with a special focus on $\boldsymbol{x}^{k}_{q}$, which represents the last query token’s representation. Without ambiguity, we will omit layer superscripts.
The output of an attention head can be expressed as follows:
$f(\boldsymbol{x}_{q})=\boldsymbol{W}_{V}\boldsymbol{X}\ \texttt{Softmax}\left(\frac{(\boldsymbol{W}_{K}\boldsymbol{X})^{T}\boldsymbol{W}_{Q}\boldsymbol{x}_{q}}{\sqrt{d_{out}}}\right),$ (4)
where $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K}$, and $\boldsymbol{W}_{V}$ are projection matrices belonging to $\mathbb{R}^{d_{out}\times d_{in}}$. These matrices are utilized to compute the attention queries, keys, and values, respectively. To simplify the notation below, we denote $\boldsymbol{W}_{Q}\boldsymbol{x}_{q}$ as $\boldsymbol{q}$ and partition $\boldsymbol{X}$ into $[\boldsymbol{X}^{\prime},\boldsymbol{x}_{q}]$.
Dai et al. (2023) describe ICL as a meta-optimization process, inspired by the previously mentioned “dual form.” In their framework, the LLM acts as the meta-optimizer, with the meta-gradient according to demonstrations generated during the forward computation and applied to the model through the attention mechanism:
$f(\boldsymbol{x}_{q})$ $\approx\boldsymbol{W}_{V}[\boldsymbol{X}^{\prime};\boldsymbol{x}_{q}](\boldsymbol{W}_{K}[\boldsymbol{X}^{\prime};\boldsymbol{x}_{q}])^{\top}\boldsymbol{q}$ (5)
$=\underbrace{\boldsymbol{W}_{V}\boldsymbol{x}_{q}(\boldsymbol{W}_{K}\boldsymbol{x}_{q})^{\top}}_{\text{Denoted as $\boldsymbol{W}_{0}$}}\boldsymbol{q}\ +$
$\ \ \underbrace{\texttt{LA}(\boldsymbol{W}_{V}\boldsymbol{X}^{\prime},\boldsymbol{W}_{K}\boldsymbol{X}^{\prime},\boldsymbol{q})}_{\text{$N$ demonstrations' effect, denoted as $\Delta\boldsymbol{W}_{N}\boldsymbol{q}$}}$
$=(\boldsymbol{W}_{0}+\Delta\boldsymbol{W}_{N})\ \boldsymbol{q}.$
In the equation above, the standard attention is approximated to a relaxed linear attention. $\boldsymbol{W}_{V}\boldsymbol{x}_{q}(\boldsymbol{W}_{K}\boldsymbol{x}_{q})^{\top}$ is denoted as $\boldsymbol{W}_{0}$ because $\boldsymbol{W}_{0}\boldsymbol{q}$ is the attention output in the 0-shot setting. According to Eq. 3, the outputs from linear attention can be interpreted as the effect of $N$ demonstrations. Consequently, this term is denoted as $\Delta\boldsymbol{W}_{N}\boldsymbol{q}$.
3 Method
[…]
3.1 The Order of ICL Examples Affects Model Outputs
From this meta-optimization perspective, we provide an explanation for the well-known issue of LLMs’ sensitivity to the order of ICL examples (Lu et al., 2022).
Considering the term $\Delta\boldsymbol{W}_{N}$ in Eq. 5, the causal attention in LLMs permits only later tokens to attend to earlier ones. This implies that altering the order of demonstration examples will correspondingly change the content of $\boldsymbol{X}^{\prime}$, thereby affecting the meta-gradients $\Delta\boldsymbol{W}_{N}$ and ultimately the resulting different obtained ICL model and its outputs. We would like to compare this meta-optimization process to optimizing a neural network: for $N$ training samples with a batch size of one, it requires $N$ back-propagation steps. During each step, the gradients are computed and iteratively update the model. The gradients for each training sample depend not only on the sample itself but also on the parameters updated by preceding samples. Therefore, the order of training examples influences the final model parameters and, consequently, its overall performance.
Despite the significant impact of the orders on performance, this explanation also applies to the observation by Lu et al. (2022) that an increase in the number of examples is correlated with greater variance in performance. This occurs because a larger set of examples raises the upper limit of potential performance. However, due to the meta-gradients being computed sequentially from individual samples, the optimization process is directed towards diverging outcomes, most of which are suboptimal.
Figure 1: (a) Standard in-context learning. (b) Batch-ICL aggregates the meta-gradients generated during individual 1-shot learning forward computations and applies them to a zero-shot forward process. (c) Multi-epoch Batch-ICL further enhances ICL performance, shown here with a 2-epoch overview.
3.2 ICL Examples Batching and Meta-Gradients Aggregation
Inspired by the analogy above, we develop an ICL inference algorithm named “Batch-ICL” which empowers LLMs to handle each ICL example and produce the corresponding meta-gradients individually, mirroring the way they are arranged in a training batch. By aggregating these meta-gradients and then applying them to a zero-shot forward process, we have observed that LLMs exhibit superior performance compared to standard ICL.
We adhere to the terminology used in Section 2. Considering a set of $N$ demonstration examples, we opt for the LLM to undertake $N$ individual 1-shot ICL learning processes, diverging from the standard single $N$-shot learning. Each sample received by LLMs is formatted as “$I_{i},l_{i},I_{q}$”, and $i$ spans the entire set $N$. At a selected layer $k$, we collect the $f_{i}(\boldsymbol{x_{q}})$ at the last position, i.e., the result of Eq. 5. Since $\boldsymbol{X}^{{}^{\prime}}$ in Eq. 5 now represents the representation of “$I_{i},l_{i}$”, as opposed to the entire set of ICL examples, we accordingly adjust the formula to: $f_{i}(\boldsymbol{x_{q}})=(\boldsymbol{W}_{0,i}+\Delta\boldsymbol{W}_{1,i})\boldsymbol{q}$.
When aggregating these hidden states $[f_{i}(\boldsymbol{x_{q}})]^{N}_{i=1}$ using the arithmetic mean, we actually aggregate meta-gradients for each 1-shot learning:
$\bar{f}(\boldsymbol{x_{q}})$ $=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\boldsymbol{x_{q}})$ (6)
$=\frac{1}{N}\sum_{i=1}^{N}(\boldsymbol{W}_{0,i}+\Delta\boldsymbol{W}_{1,i})\boldsymbol{q}$
$\approx\left(\boldsymbol{W}_{0}+\frac{1}{N}\sum_{i=1}^{N}\Delta\boldsymbol{W}_{1,i}\right)q.$
The approximation here is due to the minor variations between $\boldsymbol{W}_{0,i}$ and the actual zero-shot weight $\boldsymbol{W}_{0}$, influenced by $x_{q}$ being affected by the preceding 1-shot example.
Next, we substitute the zero-shot outputs $\boldsymbol{W}_{0}\boldsymbol{q}$ at the same layer $k$ with the aggregated $\bar{f}(\boldsymbol{x_{q}})$, to obtain the final prediction.
To determine the value for parameter $k$, taking into account the variability introduced by model size, parameters, and tasks, we employ a general approach: For any given task and LLM, we evaluate $k$ across the range of maximum layers. We then choose the $k$ that yields the best performance on the validation set. The selected $k$ remains constant during the testing phase for this LLM and task.
We name this inference algorithm as “Batch-ICL” because it is akin to computing gradients during the optimization of a neural network using a batch of inputs.
3.3 Expanding to Multiple “Epochs”
We notice that the meta-optimization perspective can be extended to encompass “multiple epochs,” which in turn further improves the performance of Batch-ICL. To clarify this concept, let’s start by distinguishing the layers with superscripts $k$ in Eq. 5:
$f^{k+1}(\boldsymbol{x_{q}}^{k})$ $=f^{k+1}(f^{k}(\boldsymbol{x_{q}}^{k})).$ (7)
Here, we simplify the feedforward layer by treating it as a linear transformation, considering it as a unit matrix for clarity purposes. When we substitute the aggregated $\bar{f}^{k}(\boldsymbol{x_{q}}^{k})$ with its individual components $f^{k}_{i}(\boldsymbol{x_{q}}^{k})$, which are the separate outputs of each 1-shot ICL at layer $k$, and then derive the attention outputs from each $k+1$ layer for further aggregation, we actually engage in a form of meta-optimization during an additional epoch. Formally, the outputs of each 1-shot ICL at layer $k+1$ now turns to:
$f_{i}^{k+1}(\bar{f}^{k}(\boldsymbol{x_{q}}^{k}))$ (8)
$=(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\frac{1}{N}\sum_{j=1}^{N}\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$
$=(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})\frac{1}{N}\sum_{j=1}^{N}(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$
$=\frac{1}{N}\sum_{j=1}^{N}(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}.$
Notice the expression $(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$, which can be understood as conducting meta-optimization for example $j$ in the first $k$ layers, followed by a similar meta-optimization for example $i$ at layer $k+1$. We term this procedure an extra “epoch” within the Batch-ICL, and it can readily be extended to multi-epoch. Figure 1(c) illustrates this 2-epoch expansion of Batch-ICL.
Essentially, the multi-epoch Batch-ICL implicitly enumerates all permutations of ICL examples, yet it achieves this more efficiently. To understand this, we formulate $f_{i}^{k+1}(\boldsymbol{x_{q}}^{k+1})$ as:
$f_{i}^{k+1}(\boldsymbol{x_{q}}^{k+1})$ $=\frac{1}{N}\sum_{j=1}^{N}(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{ji}^{k})\boldsymbol{q}$ (9)
$=(\boldsymbol{W}_{0}^{k+1}+\frac{1}{N}\sum_{j=1}^{N}\Delta\boldsymbol{W}_{ji}^{k})\boldsymbol{q},$
where $\Delta\boldsymbol{W}_{ji}^{k}$ denotes the meta-gradient produced through optimizing example $j$ in the first $k$ layers, followed by optimization of example $i$ in the $k+1$ layer. By aggregating $f_{i}^{k+1}(\boldsymbol{x_{q}}^{k+1})$ across $i=1\cdots N$, we derive:
$\bar{f}_{i}^{k+1}(\boldsymbol{x}_{\boldsymbol{q}}^{k+1})$ $=\frac{1}{N}\sum_{i=1}^{N}f_{i}^{k+1}(\boldsymbol{x}_{\boldsymbol{q}}^{k+1})$ (10)
$=(\boldsymbol{W}_{0}^{k+1}+\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\Delta\boldsymbol{W}_{ji}^{k})\boldsymbol{q},$
In this two-epoch setting, it’s evident that $N^{2}$ orders of 2-shot examples are contemplated in the summation of meta-gradients. This implicit permutation not only maintains but also, to some extent, amplifies the interaction between examples, thereby enhancing the approach, all the while effectively preserving order-agnostic properties.
4 Comparison between Batch-ICL and standard $N$-shot ICL
4.1 Experimental Setup
We conduct this preliminary study to assess Batch-ICL in comparison to the standard $N$-shot ICL. We choose widely-used open-source LLMs, Llama-2-7B and -13B Touvron et al. (2023).
Task 7B 13B
$N$-shot Batch-ICL Best $N$-shot Batch-ICL Best
SST-2 $54.9_{\hskip 1.42271pt\text{[6.41]}}$   $\textbf{58.1}_{\hskip 1.42271pt\text{[8.78]}}$ $58.3_{\hskip 1.42271pt[8.89]}$ $63.8_{\hskip 1.42271pt[15.14]}$    $\textbf{68.8}_{\hskip 1.42271pt[11.35]}$    $69.4_{\hskip 1.42271pt[11.16]}$
RTE $53.0_{\hskip 1.42271pt\text{[6.77]}}$   $\textbf{62.9}_{\hskip 1.42271pt\text{[0.85]}}$ $64.2_{\hskip 1.42271pt[0.85]}$ $78.1_{\hskip 1.42271pt[4.21]}$    $\textbf{78.5}_{\hskip 1.42271pt[1.84]}$    $81.4_{\hskip 1.42271pt[2.30]}$
QNLI $50.7_{\hskip 1.42271pt\text{[1.79]}}$   $\textbf{52.8}_{\hskip 1.42271pt\text{[1.61]}}$ $55.3_{\hskip 1.42271pt[0.83]}$ $51.1_{\hskip 1.42271pt[2.51]}$    $\textbf{56.3}_{\hskip 1.42271pt[1.48]}$    $56.8_{\hskip 1.42271pt[1.14]}$
AGNews $28.7_{\hskip 1.42271pt\text{[3.84]}}$   $\textbf{29.5}_{\hskip 1.42271pt\text{[3.02]}}$ $29.6_{\hskip 1.42271pt[3.00]}$ $30.5_{\hskip 1.42271pt[6.28]}$    $\textbf{31.8}_{\hskip 1.42271pt[3.96]}$    $31.9_{\hskip 1.42271pt[3.80]}$
MRPC $40.5_{\hskip 1.42271pt\text{[10.5]}}$   $\textbf{66.7}_{\hskip 1.42271pt\text{[0.49]}}$ $67.1_{\hskip 1.42271pt[0.17]}$ $52.4_{\hskip 1.42271pt[14.83]}$    $\textbf{56.3}_{\hskip 1.42271pt[5.93]}$    $56.4_{\hskip 1.42271pt[5.95]}$
Table 1: The comparison between Batch-ICL and standard $N$-shot ICL. We report the average score and standard deviation [$\sigma$] across 10 runs. The “Best” column reports the upper limit of Batch-ICL in which we search for the optimal $k$ for each test sample. Batch-ICL not only enhances LLMs’ ICL performance in diverse tasks but also decreases performance variability across different demonstration examples, often approaching its maximum potential in numerous scenarios. All results are passed with t-test with p-value < 0.05.
The approach is evaluated on various classification tasks, containing one sentiment detection dataset SST-2 Socher et al. (2013), two natural language inference datasets RTE Dagan et al. (2006); Wang et al. (2019) and QNLI Wang et al. (2018), a topic classification dataset AGNews Zhang et al. (2015) and one paraphrase dataset MRPC Dolan and Brockett (2005).
We evaluate performance on the test set when labels are accessible, and we sample 300 data from the validation set to derive the optimal value of $k$. In cases where test labels are not provided, we evaluate performance on the validation set and choose 300 data from the training set for obtaining $k$. In this comparison, we fix the value of $N$ at 4.
For subsequent experiments in this paper, unless specifically stated otherwise, we present the average score obtained from 10 runs, each with different sampled demonstrations. This ensures reliable conclusions, as different ICL demonstrations lead to varying performance (Zhao et al., 2021; Perez et al., 2021).
4.2 Experimental Results
We present the experimental results in Table 1. Overall, Batch-ICL demonstrates a substantial improvement over the standard $N$-shot ICL in various tasks and across different model sizes. To determine Batch-ICL’s maximum potential, we identified the optimal $k$ for each test sample and have included these findings in the “Best” column. These results show that the performance discrepancy between our chosen $k$ and the theoretical upper limit of performance is negligible. Furthermore, a larger batch size used during meta-optimization results in reduced noise and randomness in the meta-gradients. Consequently, we observe Batch-ICL exhibits markedly more stable performance across different demonstrations, as evidenced by the significantly reduced variation [^1] An exception occurs with Llama-2-7B in the SST-2 task, where Batch-ICL increases variance. This happens because, in this binary classification task, the “$N$-shot” is close to stable random guessing. in repeated experiments. The improved accuracy and reduced variance serve as initial validation for the effectiveness of Batch-ICL, and also support our explanation in Section 3.1.
5 Method Analysis
In this section, we delve into various aspects of Batch-ICL, including the time cost, the impact of shot number ($N$), aggregation layer index ($k$), the order of ICL examples, and the optimization “epochs.” This exploration aims to provide a deeper insight into both the ICL task itself and the efficacy of Batch-ICL.
5.1 Discussion on the Efficiency
We analyze the efficiency of Batch-ICL in comparison to the standard $N$-shot ICL approach. Our main focus lies on addressing the computational overhead linked to the attention mechanism in Transformer models. This part is a critical efficiency bottleneck for Transformers, showcasing a time complexity of $O(n^{2})$ when processing texts of length $n$.
Consider inputs $(I_{1},l_{1},I_{2},l_{2},\dots,I_{N},l_{N},I_{q})$, and suppose the average length of both an input and a label is $T$. In a standard $N$-shot ICL task, the total length of $N$ examples combined with the query is $(2N+1)T$, resulting in an approximate time cost of $O((2N+1)^{2}T^{2})$. In contrast, Batch-ICL deconstructs the $N$-shot scenario into $N$ separate 1-shot instances, which are subsequently integrated with a 0-shot learning. Specifically, we process $(I_{i},l_{i},I_{q})$, each with a length of $3T$, $N$ times, and $(I_{q})$ with a length of $T$ once, culminating in a total time cost of approximately $10T^{2}$.
When $N>2$, it is obvious that Batch-ICL outperforms the standard ICL in terms of efficiency. It’s important to note that, in practical applications, we execute the $N$ 1-shot learning tasks simultaneously in a batch, which effectively reduces the actual latency. Furthermore, in many situations, particularly in classification tasks, the length of the label is considerably shorter than that of the input. Therefore, the benefits of our approach are even more pronounced.
5.2 The Effect of $N$
We first delve into an exploration of the impact of $N$, which plays a crucial role in task settings. The modification of $N$ poses challenges to the robustness in various practical applications. As illustrated in Figure 2, our Batch-ICL consistently surpasses the standard $N$-shot accuracy for different values of $N$. Also, it is important to take note of the gap that exists between the performance of Batch-ICL and its upper limit referred to as “Best.” This gap is minor and diminishes as the value of $N$ increases, implying improved robustness and performance.
Figure 2: Performance dynamics across various $N$.
5.3 The Effect of $k$
Figure 3: Performance dynamics across various aggregation layer $k$.
We perform a thorough analysis of the impact of $k$. We examined the effects of $k$ by varying its value across the full range of layers, monitoring the performance dynamics as aggregated meta-gradients are implemented at various levels. For each value of $k$, we carried out experiments using the Llama-2-13B model on the SST-2 dataset (Socher et al., 2013), with $N$ values ranging from 2 to 6.
Figure 3 demonstrates that with a small $k$, Batch-ICL does not work. It suggests that the initial, or shallow, layers of LLMs play a crucial role in establishing semantic foundations for the subsequent, deeper layers. This observation aligns with previous studies (Wang et al., 2023; Todd et al., 2023; Hendel et al., 2023).
As the forward computation goes on, we observe an abrupt increase in performance. This improvement plateaus rapidly and does not diminish thereafter. This pattern differs from the findings reported in (Hendel et al., 2023; Todd et al., 2023), where researchers found that a deeper representation of an ICL task diminishes the task-related information it contains. Specifically, when the representation is extracted and used in a zero-shot forward computation, the accuracy tends to converge to zero as the layer depth increases. The discrepancy between ours and theirs findings might be due to subtle differences in input configuration. In our approach, each 1-shot forward computation uses a true query. In contrast, previous studies used a pseudo query, which can lead to significant deviations in deeper layers because most information gathers at the last position (Wang et al., 2023).
Figure 4: Comparing Batch-ICL and standard ICL with various example orders, including the “Best”, “Worst” and “Average” of all permutations.
Model $N$ 2 3 4 5
Llama-2-7B $60.0$ $90.0$ $65.8$ $63.4$
Llama-2-13B $80.0$ $86.7$ $70.4$ $60.8$
Table 2: The percentage of example orders that Batch-ICL outperforms among all possible example permutations.
$N$ 7B 13B
1-epoch 2-epoch 3-epoch 4-epoch 1-epoch 2-epoch 3-epoch 4-epoch
2 $57.9$   $58.4$ $58.8$ 59.0    $69.8$    70.0 70.0 70.0
3 $60.3$   $63.4$ $64.8$ 67.0    $71.2$    71.6 71.6 $71.3$
4 $58.3$   $60.2$ $60.7$ 62.2    $69.4$    70.1 $70.0$ $69.7$
5 $57.5$   $59.5$ $60.1$ 60.4    $71.7$    71.8 $71.7$ $71.3$
6 $61.0$   $62.7$ $63.2$ 64.6    $76.2$    76.3 76.3 $75.8$
Table 3: Multiple epochs prove to enhance accuracy, irrespective of model size and shots. Note that, in this study, we do not determine $k$ based on the validation set but rather employ the optimal value for each individual sample.
5.4 The Effect of ICL Example Order
We now delve into examining the order of ICL examples and its impact on performance.
We meticulously examine all $N!$ possible permutations of ICL examples in a $N$-shot SST-2 (Socher et al., 2013) task. In Figure 4, the “Average” result is the average accuracy of all permutations. “Best Order” denotes the highest accuracy, contrasting with the “Worst Order” which represents the lowest.
It is clear that, regardless of the model’s size, the performance under the worst order is close to a random guess. In stark contrast, the best order significantly surpasses the average performance. This emphasizes the critical role of sequence organization in in-context learning (Lu et al., 2022).
In situations where only a few examples are available (e.g., $N$=2 or 3), Batch-ICL not only exceeds the average performance but also surpasses the best order. This underscores the strength of our proposed method in data-limited scenarios. As the number of examples increases, the number of permutations rises, enhancing the likelihood of achieving high accuracy; consequently, the best-order performance improves markedly. Nevertheless, Batch-ICL consistently demonstrates superior performance compared to the average.
Additionally, we performed a statistical analysis on the percentage of permutations that Batch-ICL outperforms. It is obvious that Batch-ICL exceeds the majority of permutations in every comparison. This is particularly noticeable for smaller values of $N$ or with larger models. Overall, this study confirms our motivation, demonstrating that Batch-ICL effectively alleviates concerns related to ICL example orders, often leading to a satisfactory solution.
5.5 The Effect of Epochs
In Table 3, we present the results of adding more epochs to Batch-ICL on SST-2 (Socher et al., 2013). This table illustrates the model’s performance across various numbers of $N$ and a range of epochs. Our findings indicate that, regardless of the value of $N$ or the size of the model, there is an improvement in the model’s performance over several epochs, culminating in a plateau.
We observed a more pronounced effect in the 7B model compared to the 13B model. The 7B model consistently shows enhanced performance with an increasing number of epochs, peaking at 4 epochs. On the other hand, the 13B model reaches the plateau earlier. This could be due to the larger capacity of the 13B model, which enables it to more effectively capture linguistic subtleties and contextual details with fewer implicit interactions among ICL examples (see Eq.10).
6 Related Works
The theoretical foundation of our work, including the dual form between linear attention and linear layers optimized by gradient descent, and the interpretation of In-Context Learning (ICL) from a meta-optimization perspective, is grounded in various studies  (Aizerman et al., 1964; Irie et al., 2022; Dai et al., 2023; Ren and Liu, 2023). Additionally, our research is related to the findings of Hendel et al. (2023) and Todd et al. (2023). These works discovered that in few-shot learning tasks, the representation in deep layers carries the task information. Hendel et al. (2023) focused on identifying and aggregating the outputs of induction heads (Olsson et al., 2022). Their work emphasized the content analysis of aggregated hidden states. In contrast, our approach does not differentiate between attention heads, leading to a simpler yet effective method for enhancing the ICL performance. When compared to Hendel et al. (2023), our method demonstrates greater efficacy and surpasses the original few-shot learning performance by using zero-shot learning, an achievement not realized by Hendel et al. (2023). Theoretically, our work offers an explanation for the “training set compression” proposed by Hendel et al. (2023) and elucidates why averaged attention activation, as investigated in (Todd et al., 2023), conveys ICL task information.
7 Conclusion
In this paper, from a meta-optimization perspective, we provide an explanation for ICL example orders’ impact on performance. Building on our insights, we introduce Batch-ICL, an efficient and effective algorithm for ICL inference. Batch-ICL processes ICL examples in batches, aggregating their meta-gradients. These aggregated meta-gradients are then applied to a zero-shot forward computation for final predictions. Owing to its batch processing approach, Batch-ICL remains agnostic to the order of ICL examples, surpassing the average performance of all order permutations across various tasks. We further enhance Batch-ICL by developing multi-epoch variants that implicitly enumerate permutations of ICL examples. This enhancement fosters better interaction between inputs, thereby further improving the method.
Acknowledgement
This work was supported by Ant Group.